OpenStudy (anonymous):

Saira Weights 25 pounds more than Umbar. If together they weight 205 pounds, what is the weight of Saira ?

OpenStudy (anonymous):

let Umbar weighs X pound, then Saira weighs X+25 pounds and together they weigh X+X+25 = 205 pounds so 2X+25=25=>2X=180 so X=90 pounds, recall Saira was X+25 so 90+25=115 pounds

OpenStudy (anonymous):

i meant 2X+25=205 there

OpenStudy (anonymous):

Thanks ^^ That was helpful.

OpenStudy (anonymous):

ur welcome